# Confining quantum field theories

### v1

@@ Abstract
It is widely believed, and axiomatically postulated in mathematical quantum field theory, that the vacuum is a unique vector state (up to a phase factor). The recent solution of the quantum Yang-Mills theory of the strong interaction revealed the presence of two vacua and a mixed quantum state. The second, confining vacuum, is an eigenstate of an auxiliary field, with a non-zero eigenvalue, as opposed to the zero eigenstate of the perturbative vacuum.
Here, I show that this non-trivial vacuum structure is characteristic of confining quantum field theories, and implies confinement, in the sense that vacuum expectation values between states separated at large, space-like distances, tend to zero, whereas in ordinary quantum theories with a unique vacuum, they are known to satisfy the cluster decomposition principle, and at large separations tend to free, asymptotic states.
I conclude with a discussion and some tentative comments regarding the general status, problems, and interpretation of quantum field theory for the four known interactions.
@@ 1 Introduction
Also, in the case of Yang-Mills theory, there is a vacuum energy density difference between the two vacua that depends on $\mu^{2}$, the non-zero eigenvalue of the non-trivial operator, $\lambda^{2}$. I explain how the approach to the zero limit of the aforementioned correlations is related to this energy, but I also note that, in the general case, the non-trivial operators and vacua do not have to correspond to an energy difference. Confinement is a general characteristic of the non-trivial structure with multiple vacua, and all states in such theories are confined.
In theories where there is such an energy difference between different vacua, however, as explained in [3], there is no Lagrangian description of the theory; if such a gauge- and Lorentz-invariant description existed, one would be able to connect the two vacua using operators of the Lagrangian, and thus lower the energy, which is not possible, by the definition of a vacuum. In general confining theories, with a mixed vacuum state, the absence of asymptotic states also implies the absence of an S-matrix, and makes traditional interpretations [4, 5, 6] even more difficult.
In Sec. 2, I review the solution of Yang-Mills theory. In Sec. 3, I show the general confining property of such theories, and in Sec. 4, I make some general, more or less tentative comments, regarding the general status of quantum field theory and the other physical interactions.
@@ 2 The Yang-Mills theory
Thus, the inverted potential, $-U$, has a local minimum at $\lambda=0$, and a global maximum at $\lambda^{2}=\mu^{2}$, but because of Gauss’s law and the presence of gauge kinetic terms, the analysis of [2] demonstrated that they are both stable vacua, and in [3] it was shown that they should be interpreted as eigenstates of the auxiliary, gauge-invariant operator, $\lambda^{2}$.
This is a new mechanism of scale generation. Although the combinatorics of the Coleman-Weinberg effective potential are well known, arising from the infrared singularities of the massless, self-interacting, gauge fields, here there is no spontaneous symmetry breaking, $\lambda$ is an auxiliary field, without a kinetic term, and without any additional degrees of freedom, and $\mu^{2}$ is an eigenvalue, not a vacuum expectation value from spontaneous symmetry breaking.
@@
A rich structure of the proposed Feynman rules and effective action was already derived at previous works, including the confining interaction, soliton solutions connecting the two vacua, as well as the possible solution of the strong-CP problem, a first-order phase transition at high temperatures, and chiral symmetry breaking after the inclusion of fermions. These will be further investigated in future works. Here, as described before, I will focus more on the vacuum structure of Yang-Mills theory, and its possible generalizations.
First, it should be noted that there is an energy density difference between the two vacua: the vacuum energy density of $\Omega_{0}$ is zero, and the vacuum energy density of $\Omega_{\mu}$ is positive (equal to $-U(\mu^{2})$). The canonical formalism in [2] derives the energy (Hamiltonian) as
@@
However, since both vacua are stable, there is no well-defined Lagrangian or energy-momentum tensor that connects the two vacua. If there was, it would be possible to perturb one vacuum, using operator configurations from the Lagrangian, and lower the energy of one vacuum, which is not possible. The effective action of (1) is given at a particular Lorentz frame, in order to express the Gauss’s law constraint and derive the force between two static sources. The physics at each vacuum, however, as well as overall, is Lorentz invariant. At the vacua there is a well-defined energy-momentum tensor, $T^{\mu\nu}=T^{\mu\nu}_{\rm YM}-g^{\mu\nu}U$, where $T^{\mu\nu}_{\rm YM}$ is the usual energy-momentum tensor for perturbative Yang-Mills, and $g^{\mu\nu}$ is the Lorentz metric.
The canonical analysis also showed that the auxiliary operator, $\lambda^{2}$, is also a non-trivial operator that commutes with all other operators of the theory, without being identically zero, or a multiple of the identity. The associated non-trivial, non-unique vacuum structure, and the decomposition into different, distinct vacua, that are eigenstates of this operator, was shown to correspond to fundamental constructions of operator algebras, namely the Gelfand-Naimark-Segal (GNS) theorem.
Also, the breakdown of the cluster decomposition principle, which is equivalent to the uniqueness of the vacuum, was shown to be related to the above structure, as well as the confining properties of the theory. In the next Section, I show that confinement is characteristic of every theory with a similar, non-trivial, vacuum structure, with different vacua that are eigenstates of a non-trivial operator that commutes with every other operator of the theory.
@@ 3 Vacuum state and confinement
Pure states of the physical system are described by normalized vectors, $|\Psi>$, of a Hilbert space, and probabilities for physical observations at such pure states are given by inner products, $<\Psi|Q|\Psi>$, whereas mixed states are described by density matrices.
@@
$<\Psi_{2}|Q(f)|\Psi_{1}>=0\,\,\,\,\,{\rm if}\,\,\,\,(E_{1}+\Delta)\bigcap E_{2}=\emptyset.$ (7)
The vacuum of the theory, a vacuum state, $|\Omega>$, is defined by the property that $<\Omega|Q^{*}(f)Q(f)|\Omega>=0$, for any operator, $Q(f)$, that lowers the energy (the support of $\tilde{f}(p)$ is not in the positive cone, $p^{0}>0$, $p^{2}>0$).
@@
$<\Omega|Q_{1}(x_{1})Q_{2}(x_{2})|\Omega>\,\,\,\rightarrow\,\,<\Omega|Q_{1}(x_{1})|\Omega><\Omega|Q_{2}(x_{2})|\Omega>.$ (9)
Obviously, (9) is the cluster decomposition property, that expresses the fact that two interacting states of a system, when separated at large distances become non-interacting. This is true for theories like quantum electrodynamics, but cannot possibly be expected to hold for confining theories, like the strong interaction.
First, one notes that any operator, $Q$, can be decomposed in three parts, as $Q=Q^{-}+Q^{+}+Q^{0}$, that satisfy $Q^{-}\Omega_{\mu}=0$, $Q^{+*}\Omega_{\mu}=0$, and with the spectral support of $Q^{0}\Omega_{\mu}$ being a bounded subset of the forward and backward light cones of the energy-momentum space with $|p^{0}|\leq c$, for any arbitrary constant, $c$.
@@
Then
$<\Omega_{\mu}|Q_{1}(x_{1})Q_{2}(x_{2})|\Omega_{\mu}>=$
$<\Omega_{\mu}|(Q^{-}_{1}(x_{1})+Q^{0}_{1}(x_{1}))(Q^{+}_{2}(x_{2})+Q^{0}_{2}(x_{2}))|\Omega_{\mu}>,$ (11)
and for large $R$ one can commute $Q^{+}_{2}$ to the left and $Q^{-}_{1}$ to the right, where they give zero. The approach to zero from the commutators will be discussed later.
The remaining term is $<\Omega_{\mu}|Q^{0}_{1}(x_{1})Q^{0}_{2}(x_{2})|\Omega_{\mu}>$, and the spectral supports, $E_{1}$, $E_{2}$, of the vectors $|\Psi_{1}>=Q^{0}_{1}|\Omega_{\mu}>$, $|\Psi_{2}>=Q^{0}_{2}|\Omega_{\mu}>$ are as described above, bounded subsets of the energy-momentum light-cones.
Since $|\Omega_{\mu}>$ is an eigenvector of the operator $\lambda^{2}$, with a non-zero eigenvalue, and this operator commutes with any $Q$, this remaining term is also proportional to $<\Omega_{\mu}|Q^{0}_{1}(x_{1})\lambda^{2}(f)Q^{0}_{2}(x_{2})|\Omega_{\mu}>$.
Then, for spectral supports, $E_{1,2}$, that are bounded subsets of the light cones, it is easy to see that one can pick a function, $f$, with spectral support, $\Delta$, such that $(E_{1}+\Delta)\bigcap E_{2}=\emptyset$, and use (7), thus giving the final zero result in (8).
It is also interesting to notice the rough resemblance of this proof to the Feynman diagram of Fig. 5 of [2], which was used in order to derive the confining interaction.
The neglected terms from the commutators at large spacelike distances decrease to zero, as in ordinary field theory, at least at the order of $1/R$. In the case of Yang-Mills theory, with the confining vacuum having an energy density of order $\mu^{4}$, and for states localized in a physical volume of order $V$, terms of order $e^{-mR}/R$ are expected, with $m\sim\mu^{4}\,V$. For pure Yang-Mills, however, without fermion fields or other scales, $V$ is arbitrary, and a Coulomb (or Luscher-type) term cannot be excluded. Strictly speaking, in the sense of conventional field theories, there is no “mass gap”, at least not in this proof ($c$ is arbitrary) but there is an energy density difference for all states on the confining vacuum, as described before.
The important point of this work, however, is the result (8), which is a general confining statement, for any theory that has a similar structure of non-trivial vacua, and commuting non-trivial associated operators $\lambda$ (there may be more than one) with vacuum eigenstates. It is to be contrasted with the standard, unwarranted assumption (9) of conventional quantum field theories, that cannot possibly hold in confining situations.
In other confining theories, these related operators $\lambda$ do not all have to be related to an energy density difference. In some theories, therefore, unlike the strong interaction, there may exist a Lagrangian formalism and still have confinement; these and other problems of interpretation are further discussed in the next Section.
Finally, it is obvious that a similar argument holds for any number of spacelike separated operators, $Q$, as in ordinary quantum field theory.
4 Quantum field theories, old and new

### v2

@@ Abstract
It is widely believed, and axiomatically postulated in mathematical quantum field theory, that the vacuum is a unique vector state. The recent solution of the quantum Yang-Mills theory of the strong interaction revealed the presence of two vacua and a mixed quantum state. The second, confining vacuum, is an eigenstate of an auxiliary field, with a non-zero eigenvalue, as opposed to the zero eigenstate of the perturbative vacuum, and provides a new mechanism of scale generation.
I show that this non-trivial vacuum structure implies confinement, in the sense that vacuum expectation values between states separated at large, space-like distances, tend to zero, whereas in ordinary quantum theories with a unique vacuum, they are known to satisfy the cluster decomposition principle, and tend to free, asymptotic states, at large separations.
In a confined state, the correlation functions are zero at spacelike distances larger than the scale of the theory. Accordingly, they can be non-zero only along a timelike worldline (with an associated spacelike width).
The theory is by construction unitary and Lorentz invariant, but the different vacua give a direct sum decomposition. Implications on determinism and causality, and generalizations of the confinement mechanism for theories with other symmetries and interactions are discussed.
I argue that confinement, in the generalized sense, is a necessary (certainly not sufficient) condition for proposed theories of a conscious state. Also, I discuss the relation with the measurement postulate of quantum mechanics (when the “observer” is merely a detector). I argue that confinement, in the strong interaction, is an important mechanism, similar to and possibly along with decoherence, for the measurement process.
@@ 1 Introduction
Also, in the case of Yang-Mills theory, there is a vacuum energy density difference between the two vacua that depends on $\mu^{2}$, the non-zero eigenvalue of the non-trivial operator, $\lambda^{2}$.
This is a new mechanism of scale generation. Although the combinatorics of the Coleman-Weinberg effective potential are well known, arising from the infrared singularities of the massless, self-interacting, gauge fields, here there is no spontaneous symmetry breaking. $\lambda$ is an auxiliary field, without a kinetic term, and without any additional degrees of freedom, and $\mu^{2}$ is an eigenvalue, not a vacuum expectation value from spontaneous symmetry breaking.
The theory is by construction unitary and Lorentz invariant [4, 5, 6], but the different vacua give a direct sum decomposition. I discuss implications on determinism and causality, and possible generalizations of the confinement mechanism for theories with other symmetries and interactions.
Since in a confined state the correlation functions are zero at spacelike distances larger than the scale of the theory, they can also be shown to be non-zero only along a timelike worldline (with an associated spacelike width). This leads to arguments relating confinement with theories of consciousness, but also of quantum measurement.
In Sec. 2, I review the solution of Yang-Mills theory. In Sec. 3, I describe the confining properties of the theory, and prove the main identity for the correlation functions. In Sec. 4, I discuss the general structure and properties of confining field theories, and implications for unitarity and causality. In Sec. 5, I discuss generalizations of the confining mechanism for other theories and symmetries. In Sec 6, I argue for consciousness as a generalized confined state. In Sec. 7, I discuss relations with theories of measurement, and in Sec. 8, I conclude with some general comments, regarding the general status of quantum field theory and the physical interactions.
@@ 2 The Yang-Mills theory
Thus, the inverted potential, U-U, has a local minimum at $\lambda=0$, and a global maximum at $\lambda^{2}=\mu^{2}$, but because of Gauss’s law and the presence of gauge kinetic terms, the analysis of [2] demonstrated that they are both stable vacua, and in [3] it was shown that they should be interpreted as eigenstates of the auxiliary, gauge-invariant operator, $\lambda^{2}$.
@@
This is a new mechanism of scale generation, different than spontaneous symmetry breaking.
Stable soliton solutions of the equations
$\nabla^{2}\lambda^{a}=\frac{\partial U}{\partial\lambda^{a}},$ (2)
derived from (1), were also shown to exist, with a spatial extent of the order of $R_{\rm sol}\sim 1/(\alpha_{s}\mu)$, where $\alpha_{s}=g_{c}^{2}/(4\pi)$.
There is an energy density difference between the two vacua: the vacuum energy density of $\Omega_{0}$ is zero, and the vacuum energy density of $\Omega_{\mu}$ is positive (equal to $U(\mu^{2})$). The canonical formalism in [2] derives the energy (Hamiltonian) as
@@
Both vacua are stable, and there is no well-defined Lagrangian or energy-momentum tensor that connects the two vacua. If there was, it would be possible to perturb one vacuum, using operator configurations from the Lagrangian, and lower the energy of one vacuum, which is not possible. Transitions between the vacua can only happen in the presence of finite matter density, or high temperature, via the soliton solutions that connect them. The strong-CP problem is also easily resolved in the presence of two vacua. The effective action of (1) is given at a particular Lorentz frame, in order to express the Gauss’s law constraint and derive the force between two static sources. The physics at each vacuum, however, and overall, is Lorentz invariant. At the vacua there is a well-defined energy-momentum tensor, $T^{\mu\nu}=T^{\mu\nu}_{\rm YM}-g^{\mu\nu}U$, where $T^{\mu\nu}_{\rm YM}$ is the usual energy-momentum tensor for perturbative Yang-Mills, and $g^{\mu\nu}$ is the Lorentz metric. The results can also be seen as a formal justification of the “bag model”.
The canonical analysis also showed that the auxiliary operator, $\lambda^{2}$, is a non-trivial operator that commutes with all other operators of the theory, without being identically zero, or a multiple of the identity. The associated non-trivial, non-unique vacuum structure, and the decomposition into different, distinct vacua, that are eigenstates of this operator, was shown to correspond to fundamental constructions of operator algebras, namely the Gelfand-Naimark-Segal (GNS) theorem.
Also, the breakdown of the cluster decomposition principle, which is equivalent to the uniqueness of the vacuum, was shown to be related to the above structure, as well as the confining properties of the theory. In the next Section, I show that confinement is characteristic of a theory with a similar, non-trivial, vacuum structure, with different vacua that are eigenstates of a non-trivial operator that commutes with every other operator of the theory.
@@ 3 Vacuum state and confinement
Pure states of the physical system are described by normalized vectors, $|\Psi>$, of a Hilbert space, and probabilities for physical observations at such pure states are given by inner products, $\Psi|Q|\Psi>$, whereas mixed states are described by density matrices.
@@
$\Psi_{2}|Q(f)|\Psi_{1}>=0\,\,\,\,\,{\rm if}\,\,\,\,(E_{1}+\Delta)\bigcap E_{2}=\emptyset.$ (8)
The vacuum of the theory, a vacuum state, $|\Omega>$, is defined by the property that $\Omega|Q^{*}(f)Q(f)|\Omega>=0$, for any operator, $Q(f)$, that lowers the energy (the support of $\tilde{f}(p)$ is not in the positive cone, $p^{0}>0$, $p^{2}>0$).
Once the vacuum state is defined and established, any other state of the theory can be built from it, in the sense that it can be approximated as $Q|\Omega>$, with arbitrarily large precision, where $Q$ is an operator of the theory. Namely, the vacuum is a cyclic state that defines the theory. This is either postulated or built in the GNS construction.
@@
$\Omega|Q_{1}(x_{1})Q_{2}(x_{2})|\Omega>\,\,\,\rightarrow\,\,<\Omega|Q_{1}(x_{1})|\Omega><\Omega|Q_{2}(x_{2})|\Omega>.$ (10)
Obviously, (10) is the cluster decomposition property, which expresses the fact that two interacting states of a system, when separated at large distances become non-interacting. This is true for theories like quantum electrodynamics, but cannot possibly be expected to hold for confining theories, like the strong interaction.
In fact, I will prove a stronger statement, namely that
$\Omega_{\mu}|A\,Q_{1}(x_{1})Q_{2}(x_{2})\,B|\Omega_{\mu}>\,\,\,\rightarrow\,\,0,$ (11)
at large spacelike separation $R=|\vec{x_{1}}-\vec{x_{2}}|$, for any operators $A,B$, of the theory, and then generalize it for other correlation functions of the theory.
First, one notes that any operator, $Q$, can be decomposed in three parts, as $Q=Q^{-}+Q^{+}+Q^{0}$, with the spectral supports of $Q^{+},Q^{-}$ on the forward and backward light cones respectively, and with the spectral support of $Q^{0}$ being a bounded subset of the light cones of the energy-momentum space with $|p^{0}|\leq c$, for any arbitrary positive constant, $c$.
@@
Then, with this decomposition in the operators, the expression in (11) becomes
$\Omega_{\mu}|(A^{-}+A^{0})\,(Q^{-}_{1}+Q^{0}_{1}+Q_{1}^{+})(Q^{-}_{2}+Q^{0}_{2}+Q^{+}_{2})\,(B^{+}+B^{0})|\Omega_{\mu}>.$ (13)
and, for large $R$, one can commute $Q_{1}^{+}$ to the right and $Q_{2}^{-}$ to the left in this expression. The resulting vectors have spectral support $p^{0}>-c$ (the vector at the right side) and $p^{0}<c$ (the vector at the left).
Now, $|\Omega_{\mu}>$ is an eigenvector of the operator $\lambda^{2}$, with a non-zero eigenvalue, and commutes with any operator. So, for distances $R$ that are larger than the soliton radius, $R_{\rm sol}$, and taking the constant, $c$, of the analysis to be smaller than the soliton energy, $E_{\rm sol}$, we see that we can place a soliton operator, $Q_{\rm sol}$, between 0 and $R$, between the two vectors in the expression of (11), and shift the spectra of the vectors so that they do not overlap. The soliton operator, being a solution of (2), is composed of the $\lambda$ operator and commutes with other operators. Then, using (8), we get the zero result for the expression in (11).
There is a straightforward generalization of this result to a statement for general correlation functions of the theory, namely that
$\Omega_{\mu}|T(Q_{1}(x_{1})Q_{2}(x_{2})...Q_{N}(x_{N}))|\Omega_{\mu}>=0,$ (14)
(for any $x_{i},N$) unless the $x_{i}$ lie in a timelike worldline (with a spacelike width of the order of the scale of the theory).
Implications of this statement will be discussed in the next Sections. It should be noted that its proof used the extension of the Hilbert space of the theory from ${\cal H}_{0}\oplus{\cal H}_{\mu}$, build on the asymptotic and confining vacua $\Omega_{0}$ and $\Omega_{\mu}$, to ${\cal H}_{0}\oplus{\cal H}_{\mu}\oplus{\cal H}_{\rm sol}$, that includes the soliton Hilbert space. It is possible to derive the same statement solely in ${\cal H}_{0}\oplus{\cal H}_{\mu}$, instead of placing a soliton operator between spacelike separated points, we can consider an energy increasing tube of the confining vacuum, making essentially a statement for the bag model, which expresses the fact that the strong interactions are confined at the scale of the theory for energy reasons.
4 Confining and asymptotic quantum field theories